OpenStudy (anonymous):

H10P3: An L Network if f=630.0kHz Cmatch=????? Lmatch=?????

OpenStudy (nurali):

Cmatch= 1101.75 pF Lmatch= 55.05 uH

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)

OpenStudy (nurali):

Duty Cycle 58 and check tran.

OpenStudy (anonymous):

with diode = 100 and Vin=3v

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is correct

OpenStudy (anonymous):

last question plzz H9P1: Response to a Delayed Impulse L=20.0H and C=20.26mF. A=2/π=0.64 Coulombs at time t=9.0s. B. At the initial time what is the total energy, in Joules, stored in the circuit? C. At the time just before the impulse happens t=9.0− what is the total energy, in Joules, stored in the circuit? E. At the time just before the impulse happens what is the voltage vC(9.0−), in Volts, across the capacitor? incorrect

OpenStudy (anonymous):

last question plzz H9P1: Response to a Delayed Impulse L=20.0H and C=20.26mF. A=2/π=0.64 Coulombs at time t=9.0s. B.==? C.==? E.==?

OpenStudy (anonymous):

H9P1: Response to a Delayed Impulse L=20.0H and C=20.26mF. A=2/π=0.64 Coulombs at time t=9.0s. B.==? C.==? E.==?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

32.5 incorrect 32.5 incorrect -102.05 incorrect

OpenStudy (anonymous):

No

OpenStudy (anonymous):

red marks

OpenStudy (nurali):

ok wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry to say all 3 are incorrect

OpenStudy (anonymous):

again

OpenStudy (anonymous):

no

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

Note: In this problem we have chosen numbers for the part parameters to make it easier to compute an answer :-). By the way, it is also hard to arrange zero resistance, except with superconducting materials at very low temperatures. In the circuit shown below L=20.0H and C=20.26mF. The current source puts out an impulse of area A=2/π=0.64 Coulombs at time t=9.0s. At t=0 the state is: vC(0)=0.0 and iL(0)=1.0. The equation governing the evolution of the inductor current in this circuit is d2iL(t)dt2+1LCiL(t)=ALCδ(t−9.0) What is the natural frequency, in Hertz, of this circuit? correct At the initial time what is the total energy, in Joules, stored in the circuit? incorrect At the time just before the impulse happens t=9.0− what is the total energy, in Joules, stored in the circuit? incorrect At the time just before the impulse happens what is the current iL(9.0−), in Amperes, through the inductor? correct At the time just before the impulse happens what is the voltage vC(9.0−), in Volts, across the capacitor? incorrect At the time just after the impulse happens what is the current iL(9.0+), in Amperes, through the inductor? correct At the time just after the impulse happens what is the voltage vC(9.0+), in Volts, across the capacitor? correct At the time just after the impulse happens what is the total energy, in Joules, stored in the circuit? correct

OpenStudy (nurali):

15 15 -47.1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

capcitor and inductor are in parallel

OpenStudy (anonymous):

7.5 7.5 -23.56 is also wrong

OpenStudy (nurali):

10 10

OpenStudy (anonymous):

40 40 -40*pi

OpenStudy (anonymous):

h9p1 --- e) -15.7079632679 voltage across capacitor